OpenStudy (he66666):

How does (1/y)(dy/dt) = (d/dt)ln|y| ? Can someone please help me?

hartnn (hartnn):

heard about chain rule ?

OpenStudy (he66666):

Yes, I know the chain rule and I know that (d/dx)In|y| =1/y..

hartnn (hartnn):

nopes, (d/dx)In|y| =1/y*(dy/dx)

OpenStudy (he66666):

I don't get why the (dy/dx) is there :/

hartnn (hartnn):

y is the function of x , right ? ACTUALLY d/dy (ln y)=1/y

hartnn (hartnn):

to have dx in denominator, d/dx = (d/dy) (dy/dx) hence we multiply by dy/dx

hartnn (hartnn):

got that or should i elaborate ?

OpenStudy (anonymous):

\[\frac{d (\ln(y))}{dt}=\frac{d (\ln(y))}{dy}\frac{d y}{dt}\]

OpenStudy (he66666):

@hartnn I understand your explanation but I just don't get how (1/y)(dy/dt) becomes (d/dt)ln|y| . I understand how (d/dt)ln|y| becomes (1/y)(dy/dt) but not the other way around

OpenStudy (he66666):

@henpen Yes, I get it but how does (1/y)(dy/dt) become (d/dt)ln|y| ?

OpenStudy (anonymous):

d(lny)/dy=1/y

OpenStudy (he66666):

Oh I see. So I guess you just substitute d(lny)/dy into 1/y then solve it?